# Graph Quantum Magic Squares and Free Spectrahedra

### v2

@@ 1 Introduction
This classical setting naturally generalizes to the noncommutative case, where the entries of the matrix are no longer numbers but positive elements of an algebra or, more commonly, of a $C^{*}$-algebra.
In their 2020 paper, De les Coves, Drescher, and Netzer studied quantum magic squares, whose entries are complex square matrices, and investigated their various properties, classifying them according to the structure of their entries.
@@
Their work provides the main foundation for the present paper, which we view as a continuation of their approach. Motivated by these results, we introduce a variant of quantum magic squares subject to constraints imposed by a graph $\Gamma$, where the underlying combinatorial structure is encoded by the adjacency matrix of $\Gamma$. We call these objects graph quantum magic squares (GQMS).
In analogy with [DDN20], we may ask for the graph-analogue of the Birkhoff–von Neumann question. Recall that
@@
The paper is organized as follows. In Section 2, we introduce the notation and basic background. We briefly recall the definition of classical magic squares and then review the setting used by De las Cuevas, Drescher, and Netzer to define quantum magic squares, together with their classification according to the properties of the entries. We also summarize the main tools employed in their 2020 work to disprove the Birkhoff–von Neumann theorem in the quantum case. Finally, we recall the definition of free spectrahedra following Evert and Helton’s framework [EH19], reformulating it in terms of linear matrix inequalities (LMI).
@@ 2.1 Quantum Magic Squares
In the noncommutative setting, matrix entries are elements of a unital $C^{*}$-algebra. A natural generalization of classical magic squares is to consider block matrices whose entries are positive elements summing to the unit along each row and column. Throughout, we restrict to the case of $\operatorname{Mat}_{s}(\mathbb{C})$.
To describe the “row sums” and “column sums” in this context, it is convenient to recall the standard terminology for operator-valued measurements.
@@
If, in addition, each $A_{i}$ is a projection ($A_{i}^{2}=A_{i}=A_{i}^{*}$), the family is called a projection-valued measure (PVM). Thus, PVMs form a special subclass of POVMs.
@@
A quantum magic square of size $n$ and internal dimension $s$ is a block matrix
@@
We refer to (1) as the magic relations. Equivalently, each row and each column forms a POVM.
Following [DDN20], we define the families:
@@
Setting $\mathcal{M}^{(n)}:=\bigcup_{s\in\mathbb{N}}\mathcal{M}_{s}^{(n)}$ and similarly $\mathcal{P}^{(n)}$, $\mathcal{C}^{(n)}$, we have
@@
For $n=1,2,3$ one has $\mathcal{C}^{(n)}=\mathcal{P}^{(n)}$. In particular, every $3\times 3$ quantum permutation matrix is commuting (see, e.g., [WB23]*Lemma 2.5).
The question raised in [DDN20] naturally leads to the framework of matrix convexity, which generalizes classical convexity to tuples of matrices and plays a central role in the study of noncommutative analogues of the Birkhoff–von Neumann theorem.
@@
$\sum_{i=1}^{k}V_{i}^{*}\,A^{(i)}\,V_{i}\ \in\ R_{t},$
where the conjugation acts blockwise, i.e., $\bigl(\sum_{i}V_{i}^{*}A^{(i)}V_{i}\bigr)_{kl}=\sum_{i}V_{i}^{*}A^{(i)}_{kl}V_{i}$. Equivalently, $R$ is closed under matrix-valued convex combinations.
For a subset $S\subseteq\operatorname{Mat}_{n}(\operatorname{Her}_{s}(\mathbb{C}))$, the matrix convex hull $\operatorname{mconv}(S)$ is the smallest matrix convex set containing $S$.
@@
The key tool used for proving this separation is the following criterion:
@@
The next result provides an explicit separation between classical and quantum magic squares.
@@
Thus, the matrix convex hull of quantum permutation matrices does not cover the set of quantum magic squares. This motivates the study of intermediate structures, such as graph quantum magic squares, which impose combinatorial symmetry constraints and as we will show in Section 3.3 still admit a free-spectrahedral description.
@@ 2.2 Free spectrahedra and linear matrix inequalities
The connection to free spectrahedra is particularly useful, as it allows us to translate combinatorial constraints (e.g., graph symmetries) into semidefinite inequalities.
To investigate the matrix convex structure of quantum magic squares, it is convenient to recall the framework of free spectrahedra, which provides an explicit description of matrix convex sets through linear matrix inequalities (LMIs). In this section we follow the notation and conventions of [EH19], slightly adapted to our notation.
@@
Fix a tuple $A=(A_{1},\dots,A_{g})\in(\mathrm{Sym}_{s}(\mathbb{R}))^{g}$. The associated monic linear pencil is
@@
It is convenient to denote the homogeneous linear part of the pencil by
@@
The free spectrahedron associated with $A$ is the graded set
@@
In the literature (see e.g. [EH19]), an Arveson extreme point is usually required to be irreducible, meaning that the underlying tuple admits no nontrivial common reducing subspace. Without this additional requirement, Definition 2.9 characterizes direct sums of Arveson extreme points. Since all extremal objects considered in this paper are irreducible, this distinction will not play a role in what follows.
@@
Let $\mathcal{D}$ be a compact free spectrahedron which is closed under complex conjugation. In particular, this hypothesis holds whenever the defining pencil has real coefficient matrices (since then $X\in\mathcal{D}(s)$ implies $\overline{X}\in\mathcal{D}(s)$). This will be the case for the spectrahedra arising from quantum magic squares and from the graph constraints considered later, since the corresponding coefficient matrices can be chosen real. By [EH19]*Theorem 1.3, such a set satisfies
@@
Since every absolute extreme point is, by definition, an Arveson extreme point, one has
@@ 2.3 Quantum Automorphism of Graphs
To motivate the definition of Graph Quantum Magic Squares, we recall the notion of the quantum automorphism group of a finite graph, (see e.g. [MR2146039]), following Wang’s definition of quantum permutation groups [MR1316765].
@@
In the quantum setting, instead of scalars $\{0,1\}$, we allow the entries to be elements of a unital C-algebra. A matrix is called a quantum permutation matrix if its entries are projections $p_{ij}$ such that every row and column sums to the identity.
The quantum automorphism group of $\Gamma$, denoted by $\mathrm{Aut}^{+}(\Gamma)$, is defined via its universal C-algebra $C(\mathrm{Aut^{+}}(\Gamma))$, which is generated by $n^{2}$ elements $u_{ij}$ subject to the relations making a quantum permutation matrix that commutes with the adjacency matrix:
@@
Crucially, for some graphs, this algebra is non-commutative, implying the existence of genuine quantum symmetries. A graph $\Gamma$ is said to have no quantum symmetry if the algebra is commutative, which implies $C(\mathrm{Aut}^{+}(\Gamma))=C(\mathrm{Aut}(\Gamma))$.
Conversely, if the algebra is non-commutative, the graph has genuine quantum symmetries. This distinction is fundamental to our work. Our definition of Graph Quantum Magic Squares can be seen as a convex relaxation of these quantum symmetries: instead of requiring the entries to be projections (as in $C(\mathrm{Aut}^{+}(\Gamma))$), we only require them to be positive semidefinite operators (POVMs), while maintaining the commutation condition.
In the next section we present the main developments of this paper. We introduce graph quantum magic squares and give a detailed derivation of the monic linear pencils that describe both quantum magic squares and their graph-theoretic counterparts.
@@ 3 Graph Quantum Magic Squares
In this section, we introduce a graph-compatible version of quantum magic squares, obtained by imposing commutation with the adjacency matrix of a graph $\Gamma$. The motivation is to explore how the combinatorial symmetries of a graph interact with the algebraic structure of quantum magic squares.
@@ 3.1 Definitions and preliminary observations
Let $\Gamma=(V,E)$ be a simple graph with $|V|=n$, and let $A_{\Gamma}$ denote its adjacency matrix. The external dimension $n$ of the block matrix corresponds to the number of vertices of $\Gamma$.
@@
i.e., if it commutes with the adjacency matrix of the graph tensored with the identity matrix of the internal dimension $s$. The tensor $I_{s}\otimes A_{\Gamma}$ enforces the graph structure on the internal $s\times s$ matrix entries, while leaving the external indices unchanged.
The block formulation of this commutation condition can be written as
@@
where $k\sim j$ denotes adjacency in $\Gamma$.
@@
Following the main question posed in [DDN20], we consider its graph-theoretic analogue. In this setting, we consider the matrix convex hull of the graph quantum permutation matrices, denoted by $\mathrm{mconv}(\mathcal{P}^{(\Gamma)})$, and we ask whether it covers all graph quantum magic squares:
@@
As already observed in the non-graph case [DDN20], working directly with $\mathcal{P}^{(n)}$ is often difficult, whereas $\mathcal{C}^{(n)}$ is typically much easier to analyse. For this reason, in the graph setting it is useful to understand when $\mathcal{P}^{(\Gamma)}$ and $\mathcal{C}^{(\Gamma)}$ coincide, so that one may replace $\mathcal{P}^{(\Gamma)}$ with the easier-to-handle $\mathcal{C}^{(\Gamma)}$.
Understanding whether $\mathcal{P}^{(\Gamma)}=\mathcal{C}^{(\Gamma)}$ is equivalent to asking whether the graph has no quantum symmetry, i.e., whether its quantum automorphism group reduces to the classical one. Recall that a graph $\Gamma$ has no quantum symmetry if its quantum automorphism group $\operatorname{Aut}^{+}(\Gamma)$ coincides with the classical automorphism group $\operatorname{Aut}(\Gamma)$.
Thanks to the classification results in [BB] and [SS18], many families of graphs with this property are known. For instance, the complete graphs $K_{2}$ and $K_{3}$, the cycle graphs $C_{n}$ for $n\neq 4$, and the Petersen graph all have no quantum symmetries, and therefore satisfy
@@ 3.2 A counterexample for the square graph $C_{4}$
In the general (non-graph) quantum magic square setting, [DDN20] established that the Birkhoff–von Neumann theorem does not extend to the quantum case.
@@
so the Birkhoff–von Neumann property fails for $C_{4}$ already at internal dimension $s=2$.
@@
and therefore $B\in\mathcal{M}^{(C_{4})}$. Details on the averaging construction are given in Proposition 3.4.
Next, we recall the separation criterion (Proposition 2.5): if $A\in\operatorname{mconv}(\mathcal{P}^{(4)})_{2}$, then there exists
@@
and therefore $B\notin\operatorname{mconv}(\mathcal{P}^{(C_{4})})$. This proves the claim. ∎
We now justify the averaging step used to enforce the $C_{4}$-compatibility of the counterexample.
@@
Set $\widetilde{P}_{C_{4}}:=P_{C_{4}}\otimes I_{s}$. Then $\widetilde{P}_{C_{4}}$ is unitary, so $\widetilde{P}^{*}_{C_{4}}=\widetilde{P}^{-1}_{C_{4}}$ and, for every integer $k$,
@@
We prove (ii), (iii), (i) in this order.
(ii) Commutation with $A_{C_{4}}\otimes I_{s}$.
First, we show that $\Phi(A)$ commutes with $\widetilde{P}_{C_{4}}$. Using the definition of $\Phi$ and linearity,
@@
With the change of index $m=k+1\pmod{4}$, the sum becomes
@@
Since $\Phi(A)$ is a convex combination of conjugates of $A$, it is Hermitian whenever $A$ is. Taking adjoints in $\widetilde{P}_{C_{4}}\,\Phi(A)=\Phi(A)\,\widetilde{P}$ we obtain
$\Phi(A)\,\widetilde{P}_{C_{4}}^{*}=\widetilde{P}_{C_{4}}^{*}\,\Phi(A),$
so $\Phi(A)$ commutes with both $\widetilde{P}_{C_{4}}$ and $\widetilde{P}_{C_{4}}^{*}$, hence with their sum:
@@
(iii) Entrywise formula.
We identify $A\in\operatorname{Mat}_{4}(\operatorname{Her}_{s}(\mathbb{C}))$ with its $4\times 4$ block form $A=(A_{ij})_{i,j=1}^{4}$, where $A_{ij}\in\operatorname{Her}_{s}(\mathbb{C})$. For a matrix $M$ in block form, the $(i,j)$-th block is $M_{ij}=(e_{i}^{*}\otimes I_{s})\,M\,(e_{j}\otimes I_{s})$.
@@
Since the indices are taken modulo $4$, summing over $k$ or over k-k gives the same set of terms, and we can equivalently write
$\Phi(A)_{ij}=\frac{1}{4}\sum_{k=0}^{3}A_{\,i+k,\;j+k}.$
This proves (iii).
(i) Preservation of the QMS constraints.
A quantum magic square $A$ satisfies:
@@
Conjugation by $\widetilde{P}_{C_{4}}^{k}$ simply permutes rows and columns (and hence the blocks) of $A$, so each $\widetilde{P}_{C_{4}}^{\,k}A\widetilde{P}_{C_{4}}^{-k}$ is again a QMS: the blocks remain positive semidefinite and the row/column sums stay equal to $I_{s}$.
The average $\Phi(A)$ is a convex combination of these matrices, so:
@@
Therefore $\Phi(A)$ satisfies (a), (b), (c), and is a quantum magic square. This proves (i). ∎

### v3

@@ 1 Introduction
This classical setting generalizes to the non-commutative case, where the entries of the matrix are no longer numbers but positive semidefinite elements of an algebra or, more commonly, of a $C^{*}$-algebra.
In their 2020 paper, De les Coves, Drescher, and Netzer studied quantum magic squares, whose entries are complex square matrices, classifying them according to the properties of their entries.
@@
Their work provides the main foundation for this paper, which we view as a continuation of their approach. Motivated by these results, we introduce a variant of quantum magic squares subject to constraints imposed by a graph $\Gamma$, where the underlying combinatorial structure is encoded by the adjacency matrix of $\Gamma$. We call these objects graph quantum magic squares (GQMS).
In analogy with [DDN20], we may ask for the graph-analogue of the quantum Birkhoff–von Neumann question. Recall that
@@
The paper is organized as follows. In Section 2, we introduce the notation and some background. We briefly recall the definition of classical magic squares and then review the setting used by De las Cuevas, Drescher, and Netzer to define quantum magic squares, together with their classification according to the properties of the entries. We also summarize the main tools employed in their 2020 work to prove the failure of the Birkhoff–von Neumann theorem in the quantum case. Finally, we recall the definition of free spectrahedra following Evert and Helton’s framework [EH19], reformulating it in terms of linear matrix inequalities (LMI).
@@ 2.1 Quantum Magic Squares
In the noncommutative setting, the entries of a magic square are elements of a unital $C^{*}$-algebra. A natural generalization of classical magic squares is to consider block matrices whose entries are positive elements such that the sum on each row and column equals the unit of the $C^{*}$-algebra. Throughout, we restrict to the case of $\operatorname{Mat}_{s}(\mathbb{C})$.
To describe the “row sums” and “column sums” in this context, it is convenient to recall the notion of operator-valued measure.
@@
If, in addition, each $A_{i}$ is a projection ($A_{i}^{2}=A_{i}=A_{i}^{*}$), the family is called a projection-valued measure (PVM). Thus, PVMs form a subset of POVMs.
@@
A quantum magic square of external size $n$ and internal size $s$ is a block matrix
@@
We refer to (1) as the magic relations. Equivalently, we say that each row and each column forms a POVM.
Following the convention of [DNV23], we focus on the following three families of quantum magic square according to the properties of their entries. More precisely, for any $n,s\in\mathbb{N}$, we define respectively the set of quantum magic square, the set of quantum permutation matrices and the set commuting quantum permutation matrices in the following way:
@@
Taking the union over all internal size $s$, we have the total sets $\mathcal{M}^{(n)}:=\bigcup_{s\in\mathbb{N}}\mathcal{M}_{s}^{(n)}$, $\mathcal{P}^{(n)}:=\bigcup_{s\in\mathbb{N}}\mathcal{P}_{s}^{(n)}$, $\mathcal{C}^{(n)}:=\bigcup_{s\in\mathbb{N}}\mathcal{C}_{s}^{(n)}$, and they satisfy
@@
For $n=1,2,3$ we have $\mathcal{C}^{(n)}=\mathcal{P}^{(n)}$. In particular, every $3\times 3$ quantum permutation matrix is also commuting (see, e.g., [Web23]*Lemma 2.5).
As in [DDN20] and [DNV23] we use the framework of matrix convexity, a dimension-free notion of convexity, which generalizes classical convexity. This framework plays a central role in the study of noncommutative analogues of the Birkhoff–von Neumann theorem.
@@
$\sum_{i=1}^{k}V_{i}^{*}\,A^{(i)}\,V_{i}\ \in\ R_{t}.$
Here the conjugation acts blockwise, i.e., $\bigl(\sum_{i}V_{i}^{*}A^{(i)}V_{i}\bigr)_{kl}=\sum_{i}V_{i}^{*}A^{(i)}_{kl}V_{i}$. Equivalently, we say that $R$ is closed under matrix-valued convex combinations.
For a subset $S\subseteq\operatorname{Mat}_{n}(\operatorname{Her}_{s}(\mathbb{C}))$, the matrix convex hull of $S$, denoted as $\operatorname{mconv}(S)$, is the smallest matrix convex set containing $S$.
@@
The key tool used for proving this failure is the following criterion:
@@
We recall the main result of [DDN20].
@@
Thus, the matrix convex hull of quantum permutation matrices does not recover the full set of quantum magic squares. This motivates the study of graph quantum magic squares, which impose additional constraints imposed by the structure of a graph. In Section 3.3 we show that this variant also admits a free-spectrahedral description.
@@ 2.2 Free spectrahedra and linear matrix inequalities
We now recall the notion of free spectrahedra, which provides a description of matrix convex sets through linear matrix inequalities (LMIs). In this section we follow the conventions of [EH19], adapted to our notation.
@@
Fixing a tuple $A=(A_{1},\dots,A_{g})\in(\mathrm{Sym}_{s}(\mathbb{R}))^{g}$, we define the monic linear pencil associated to $A$ as
@@
Denoting the homogeneous linear part of the pencil by
@@
Given a tuple $A\in(\operatorname{Her}_{d}(\mathbb{C}))^{g}$ and a positive integer $s$, the free spectrahedron associated with $A$ is defined as the graded set
@@
In the literature (see e.g. [EH19]), an Arveson extreme point is usually required to be irreducible, meaning that the underlying tuple admits no nontrivial common reducing subspace. Without this, Definition 2.9 characterizes direct sums of Arveson extreme points. Since all extremal objects considered in this paper are irreducible, this distinction will not play a role in what follows.
@@
Let $\mathcal{D}$ be a compact free spectrahedron which is closed under complex conjugation. In particular, this hypothesis holds whenever the defining pencil has real coefficient matrices (since then $X\in\mathcal{D}(s)$ implies $\overline{X}\in\mathcal{D}(s)$). This will be the case for the spectrahedra arising from quantum magic squares and from the graph ones considered later, since the corresponding coefficient matrices can be chosen real. By [EH19]*Theorem 1.3, such a set satisfies
@@
Since every absolute extreme point is, by definition, an Arveson extreme point, we have
@@ 2.3 Quantum Automorphism of Graphs
To motivate the definition of Graph Quantum Magic Squares, we recall the notion of the quantum automorphism group of a finite graph, (see e.g. [Ban05]), following Wang’s definition of quantum permutation groups [Wan95].
@@
In the quantum setting, instead of scalars $\{0,1\}$, the entries are elements of a unital C-algebra. A matrix is called a quantum permutation matrix if its entries are projections $p_{ij}$ such that every row and column sums to the identity.
The quantum automorphism group of $\Gamma$, denoted by $\mathrm{Aut}^{+}(\Gamma)$, is defined via its universal C-algebra $C(\mathrm{Aut^{+}}(\Gamma))$. This algebra is generated by $n^{2}$ elements $u_{ij}$ subject to the relations making a quantum permutation matrix that commutes with the adjacency matrix:
@@
For some graphs, this algebra is non-commutative, implying the existence of genuine quantum symmetries. A graph $\Gamma$ is said to have no quantum symmetry if the algebra is commutative, implying $C(\mathrm{Aut}^{+}(\Gamma))=C(\mathrm{Aut}(\Gamma))$.
On the contrary, if the algebra is non-commutative, the graph has genuine quantum symmetries. This distinction is fundamental to our work. Our definition of graph quantum magic squares can be seen as a relaxation of $C(\mathrm{Aut}^{+}(\Gamma))$: instead of requiring that the entries be projections (as in $C(\mathrm{Aut}^{+}(\Gamma))$), we only require that they are positive semi-definite operators (POVM), retaining the commutation condition.
In the next section, we introduce the notion of Graph Quantum Magic Squares and provide explicit monic linear pencils that describe both quantum magic squares and the graph ones.
@@ 3 Graph Quantum Magic Squares
In this section, we introduce a graph-variant of quantum magic squares, obtained by imposing commutation with the adjacency matrix of a graph $\Gamma$. The motivation is to explore how the combinatorial symmetries of a graph interact with the algebraic structure of quantum magic squares.
@@ 3.1 Definitions and preliminary observations
Let $\Gamma=(V,E)$ be a simple graph with $|V|=n$, and let $A_{\Gamma}$ denote its adjacency matrix. Here, the external dimension $n$ of the matrix corresponds to the number of vertices of $\Gamma$.
@@
i.e., if it commutes with the adjacency matrix of the graph tensored with the identity matrix of the internal dimension $s$.
The tensor $I_{s}\otimes A_{\Gamma}$ imposes the structure of the graph on the internal entries of the matrix $s\times s$, leaving the external indices unchanged.
This commutation condition can be written in terms of internal blocks as
@@
where $k\sim j$ means that vertices $k$ and $j$ are adjacent in $\Gamma$.
@@
Following the main question in [DDN20], we consider its graph analogue. In this setting, we consider the matrix convex hull of the graph quantum permutation matrices, denoted by $\mathrm{mconv}(\mathcal{P}^{(\Gamma)})$, and we ask whether it covers the set of all graph quantum magic squares:
@@
As already observed in the non-graph case [DDN20], working directly with $\mathcal{P}^{(n)}$ is often difficult, whereas the set $\mathcal{C}^{(n)}$ is usually much easier to handle. For this reason, in the graph setting it is useful to understand when $\mathcal{P}^{(\Gamma)}$ and $\mathcal{C}^{(\Gamma)}$ coincide, so that we may replace $\mathcal{P}^{(\Gamma)}$ with the set $\mathcal{C}^{(\Gamma)}$.
Understanding whether $\mathcal{P}^{(\Gamma)}=\mathcal{C}^{(\Gamma)}$ is equivalent to asking whether the graph has no quantum symmetry, i.e., whether its quantum automorphism group reduces to the classical one. As remarked in subsection 2.3 that a graph $\Gamma$ has no quantum symmetry if its quantum automorphism group $\operatorname{Aut}^{+}(\Gamma)=\operatorname{Aut}(\Gamma)$.
Thanks to the classification results in [BB07] and [Sch18], many families of graphs with this property are known. For instance, the complete graphs $K_{2}$ and $K_{3}$, the cycle graphs $C_{n}$ for $n\neq 4$, and the Petersen graph all have no quantum symmetries, and therefore satisfy
@@ 3.2 A counterexample for the square graph $C_{4}$
In the general (non-graph) quantum magic square setting, [DDN20] showed that the Birkhoff–von Neumann theorem does not hold in the quantum case.
@@
so the Birkhoff–von Neumann theorem fails for $C_{4}$ already at internal dimension $s=2$.
@@
and therefore $B\in\mathcal{M}^{(C_{4})}$.
Proposition 3.4 gives more details about the averaging construction.
Now, we recall the necessary condition for a matrix $A$ to be in the matrix convex hull of the set of quantum permutation matrices (Proposition 2.5). If $A\in\operatorname{mconv}(\mathcal{P}^{(4)})_{2}$, then there exists
@@
and therefore $B\notin\operatorname{mconv}(\mathcal{P}^{(C_{4})})$. ∎
Next proposition justify the averaging used in Theorem 3.3.
@@
Let us set $\widetilde{P}_{C_{4}}:=P_{C_{4}}\otimes I_{s}$. Then $\widetilde{P}_{C_{4}}$ is unitary, so $\widetilde{P}^{*}_{C_{4}}=\widetilde{P}^{-1}_{C_{4}}$ and, for every $k\in\mathbb{Z}$,
@@
(ii) We start showing that $\Phi(A)$ commutes with $\widetilde{P}_{C_{4}}$. By definition of $\Phi$ and linearity we have
@@
Changing the index $m=k+1\pmod{4}$, the sum becomes
@@
Since $\Phi(A)$ is a convex combination of conjugates of $A$, if $A$ is Hermitian $\Phi(A)$ is too. Taking adjoints in (4), we get
$\Phi(A)\,\widetilde{P}_{C_{4}}^{*}=\widetilde{P}_{C_{4}}^{*}\,\Phi(A).$
So $\Phi(A)$ commutes with both $\widetilde{P}_{C_{4}}$ and $\widetilde{P}_{C_{4}}^{*}$, and so does their sum:
@@
(iii) For a block matrix $M$, the $(i,j)$-th block is $M_{ij}=(e_{i}^{*}\otimes I_{s})\,M\,(e_{j}\otimes I_{s})$.
@@
Since the indices are taken modulo $4$, summing over $k$ is equivalent to sum over k-k, and then we can write
$\Phi(A)_{ij}=\frac{1}{4}\sum_{k=0}^{3}A_{\,i+k,\;j+k}$
proving (iii).
(i) Recall that a quantum magic square $A$ satisfies:
@@
Conjugation by $\widetilde{P}_{C_{4}}^{k}$ just permutes rows and columns (and hence the blocks) of $A$, so each $\widetilde{P}_{C_{4}}^{\,k}A\widetilde{P}_{C_{4}}^{-k}$ is again a QMS: the blocks remain positive semidefinite and the row/column sums stay equal to $I_{s}$.
$\Phi(A)$ is a convex combination of these matrices, so:
@@
Therefore, since $\Phi(A)$ satisfies (a), (b), (c), it is a quantum magic square and this proves (i). ∎